a 48 year old man is two times as old as his son. How old was the father when he was three times as old as his son?
Answer by stanbon(75887)

a 48 year old man is two times as old as his son. How old was the father when he was three times as old as his son?
--------------
father = 48
son = 24
-------------------------------
Equation:
(48+x) = 3(24+x)
48 + x = 72 + 3x
2x = -24
x = -12
----
Father was 48-12 = 36
Son was 24-12 = 12
=============================
Cheers,
Stan H.